OpenStudy (anonymous):

Solve for x. If you have multiple solutions, separate them with a comma. (3/x+5)-(5/x^2+6x+5)=(2/x+1). I got the answer of 14 but its wrong. Where is my simple stupid math mistake?

OpenStudy (phi):

I assume you know that x^2+6x+5= (x+5)(x+1) I would multiply all terms both sides by (x+5)(x+1) (and cancel) to clear the denominator. Then solve for x

OpenStudy (anonymous):

Right and I did that and I got 3x+1-5=2x+10...right?

OpenStudy (phi):

isn't the first term 3*(x+1) ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

then how did you get 3x+1?

OpenStudy (anonymous):

Since you are cancelling out (x+5) of the denominator that leaves you with (x+1) of the common denominator so it would be 3*(x+1)= 3x+1

OpenStudy (phi):

when you distribute the 3, you multiply each term inside the parens by 3

OpenStudy (phi):

you did 2(x+5) correctly: 2x+10

OpenStudy (anonymous):

oh.....there is my stupid math mistake. Thanks for your help!